### v2

# A discrete formulation for three-dimensional winding number

Abstract
For a smooth map $g\colon X\to U(N)$, where $X$ is a three-dimensional, oriented, and closed manifold, the winding number or the map’s degree is defined by $W_{3}=\frac{1}{24\pi^{2}}\int_{X}\mathrm{Tr}\left[(g^{-1}dg)^{3}\right]$. We introduce a method to compute $W_{3}$ using a discrete approximation of $X$ so that the result is manifestly quantized.
preprint: YITP-24-14
Introduction— Consider a three-dimensional closed and oriented manifold $X$. For a smooth map $g:X\to U(N)$, with $U(N)$ representing the group of $N\times N$ unitary matrices, the winding number, an integer value, is defined by the following expression:
$W_{3}[g]$ $=\frac{1}{2\pi}\int_{X}H\in\mathbb{Z},$ (1)
$H$ $=\frac{1}{12\pi}\mathrm{Tr}\left[(g^{-1}\mathrm{d}g)^{3}\right].$ (2)
The winding number $W_{3}[g]$ is pivotal in various branches of physics, including topological band theory, where it acts as the topological invariant for three-dimensional superconductors with time-reversal symmetry [1, 2], and in non-Abelian (lattice) gauge theory, appears in instanton number calculations [3, 4]. Often in these applications, the function $g$ is defined only on a finite set of lattice points for numerical analysis. Therefore, an efficient numerical formulation with lattice approximation of manifolds is an important issue.
For the first Chern number $ch_{1}$ of a line bundle with connection, a sort of two-dimensional counterpart of the winding number, a well-established discrete formulation with evident quantization exists [5, 6]. Furthermore, discrete line bundles over finite simplicial complexes have been explored, especially concerning applications in computer graphics [7]. This paper develops a method for evaluating $W_{3}[g]$ via a discretized approximation of $X$, ensuring the result remains manifestly quantized, provided the approximation of $X$ is sufficiently refined relative to the spatial variation’s length scale.
Figure 1: (a) Illustration of a $\theta$-gap. (b) Projection onto the eigenspace between two $\theta$-gaps. (c) Smearing eigenvalues over a finite set of vertices. In each panel, the unit circle in the complex plane is depicted, with blue arcs representing the spectrum of the $U(N)$-valued matrix $g(x)$ within a local region. Red lines mark the locations of $\theta$-gaps.
Formulation— Given that $H$ is a closed three-form, it is locally exact, meaning that for a local patch, there exists a two-form $B$ such that $H=dB$. To construct $B$ explicitly, we introduce a gap condition for elements of $U(N)$. A matrix $g\in U(N)$ exhibits a $\theta$-gap if none of its eigenvalues are $e^{i\theta}$ for a given real number $\theta\in[0,2\pi)$ [8], as illustrated in Fig. 1 (a). Furthermore, we define $\log_{\theta}z$ for nonzero complex number $z\in\mathbb{C}^{\times}$ as
$\log_{\theta}z=\log|z|+i\arg z,\quad\theta\leq\arg z<\theta+2\pi.$ (3)
For two distinct $\theta$-gaps, $\theta_{1}$ and $\theta_{2}$, the following relation holds:
$\log_{\theta_{1}}z-\log_{\theta_{2}}z$ =$=\begin{cases}2\pi i\times\mathrm{sgn}(\theta_{1}-\theta_{2})&(\min(\theta_{1},\theta_{2})<\arg z<\max(\theta_{1},\theta_{2})),\\ 0&(\mathrm{otherwise}).\end{cases}$ (4)
Consider $U\subset X$ as a three-dimensional subspace where $g(x)$ maintains a $\theta$-gap for $x\in U$. Let $\gamma(x)=(u_{1}(x),\dots,u_{N}(x))\in U(N)$ be a unitary matrix diagonalizing $g(x)$, i.e., $g(x)=\gamma(x)\Lambda(x)\gamma(x)^{\dagger}$ with $\Lambda(x)=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$, where $\lambda_{n}\in U(1)$ for $n=1,\dots,N$. The exact form $B_{\theta}$ is given by [9]
$B_{\theta}$ $=Q+R_{\theta},$ (5)
$Q$ $=\frac{1}{4\pi}\mathrm{Tr}[\gamma^{-1}d\gamma\Lambda\gamma^{-1}d\gamma\Lambda^{-1}],$ (6)
$R_{\theta}$ $=\frac{1}{2\pi}\mathrm{Tr}[\log_{\theta}\Lambda(\gamma^{-1}d\gamma)^{2}].$ (7)
Note that while $Q$ is independent of $\theta$, but $R_{\theta}$ does. It is evident that $X$ can be covered by $\{U_{i}\}_{i}$ such that in each patch $U$, $g(x)$ exhibits a specific $\theta$-gap $\theta_{i}$.
The unitary matrix $\gamma$ is not unique due to the transformation $\gamma\mapsto\gamma W$, where $W\in U(N)$ commutes with $\Lambda$, satisfying $W\Lambda W^{-1}=\Lambda$. This ambiguity, however, does not affect $Q$ and $R_{\theta}$. To illustrate, consider the $N$ eigenvalues divided into groups of $|I|$ degenerate ones, each with eigenvalue $\lambda_{I}$ so that $\Lambda=\bigoplus_{I}\lambda_{I}{\bf 1}_{|I|}$. Introduce a block matrix notation $A_{IJ}=(u_{i}^{\dagger}du_{j})_{i\in I,j\in J}$. The transformation matrix $W$ is expressed as $W=\bigoplus_{I}W_{I}$ as well, with $W_{I}\in U(|I|)$, modifying $A_{IJ}$ to $A_{IJ}=W_{I}^{\dagger}A_{IJ}W_{J}+\delta_{IJ}W_{I}^{-1}dW_{I}$. Consequently, $Q$ and $R_{\theta}$ can be represented as:
$Q$ $=\frac{1}{4\pi}\sum_{I,J;I\neq J}\mathrm{Tr}_{I}[A_{IJ}A_{JI}]\lambda_{J}\lambda_{I}^{-1},$ (8)
$R_{\theta}$ $=\frac{1}{2\pi}\sum_{I,J;I\neq J}\mathrm{Tr}_{I}[A_{IJ}A_{JI}]\log_{\theta}\lambda_{I},$ (9)
where $\mathrm{Tr}_{I}$ denotes the trace over indices $i\in I$. In the summation $\sum_{I,J}$, terms with $I=J$ can be excluded due to $\mathrm{Tr}_{I}[(A_{II})^{2}]=0$. This demonstrates the invariance of $Q$ and $R_{\theta}$ under the transformation $\gamma\mapsto\gamma W$.
Another noteworthy aspect is that the difference in $B_{\theta}$ between two $\theta$-gaps is a total derivative. For $0\leq\theta_{1},\theta_{2}<2\pi$, and using (4), it follows that:
$B_{\theta_{1}}-B_{\theta_{2}}=d\alpha_{\theta_{1},\theta_{2}},$ (10)
where
$\alpha_{\theta_{1},\theta_{2}}$ $=-i\ \mathrm{sgn}(\theta_{1}-\theta_{2})\mathrm{Tr}[P_{\theta_{1},\theta_{2}}\gamma^{-1}d\gamma]$
$=-i\ \mathrm{sgn}(\theta_{1}-\theta_{2})$
×$\quad\times\sum_{n;\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})}u_{n}^{\dagger}du_{n},$ (11)
with
$P_{\theta_{1},\theta_{2}}=\sum_{n;\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})}u_{n}u_{n}^{\dagger},$ (12)
the orthogonal projection onto the eigenspace for eigenvalues that fulfill $\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})$. In cases where $\theta_{1}=\theta_{2}$ or no eigenvectors meet the condition $\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})$, $\alpha_{\theta_{1},\theta_{2}}$ is simply null.
Now, we express the winding number $W_{3}[g]$ as a sum of line integrals, utilizing a cubic decomposition $L$ of the manifold $X$. (Any simplicial decomposition is equally valid.) Within each cube $c$ of the lattice $L$, we select $\theta_{c}\in[0,2\pi)$ such that $g_{x}$ for $x\in c$ exhibits a $\theta$-gap of $\theta_{c}$. Thus, $W_{3}[g]$ can be reformulated as a sum of integrals over all plaquettes:
[…]
$=\frac{1}{2\pi}\sum_{c}\int_{\partial c}B_{\theta_{c}}$
$=\frac{1}{2\pi}\sum_{p}\int_{p}(B_{\theta_{p}^{-}}-B_{\theta_{p}^{+}}).$ (13)
Here, $\sum_{p}$ runs over all plaquettes $p$ in the lattice $L$, with each $p$ being oriented. The gap parameters $\theta_{p}^{+}$ and $\theta_{p}^{-}$ correspond to the cubes adjacent to plaquette $p$, in directions parallel and antiparallel to $p$’s normal vector, respectively, as depicted in Fig. 2 (a). This formulation further simplifies to a sum of line integrals:
[…]
$=\frac{1}{2\pi}\sum_{p}\oint_{\partial p}\alpha_{\theta_{p}^{-},\theta_{p}^{+}}.$ (14)
[…]
Figure 2: (a) A plaquette $p$ within the cubic lattice, showing $\theta_{p}^{+}$ and $\theta_{p}^{-}$ as the $\theta$-gaps of cubes adjacent to $p$, aligned parallel and anti-parallel to $p$’s normal vector, respectively. The vertices $v_{0},v_{1},v_{2},$ and $v_{3}$ are sequentially labeled around the perimeter of plaquette $p$. (b) An edge $v_{a}v_{b}$ of the cubic lattice, illustrating the $\theta$-gaps of cubes adjacent to the edge $v_{a}v_{b}$.
We claim that the winding number as expressed in (14) can be calculated solely using the diagonalizing matrices $\gamma(v)$ at the vertices $v$ of lattice $L$. Diagonalizing $g(v)$ at vertices $v\in L$ yields the eigenvector and eigenvalue pairs $\{u_{n}(v),\lambda_{n}(v)\}_{n=1,\dots,N}$ for each vertex $v$. The ordering of eigenvectors $u_{n}(v)$ is such that the angles of eigenvalues ascend, satisfying $0\leq\lambda_{1}(v)\leq\cdots\leq\lambda_{N}(v)<2\pi$. (Note that eigenvalues $\lambda$ near 00 may reorder significantly under minor perturbations, yet this does not contribute to the discrete formula below.) The gap parameter $\theta_{c}$ for each cube $c$ is determined as follows: From the eight vertices of cube $c$, denoted as $v\in c$, we derive $8N$ eigenvalues $\{\lambda_{n}(v)\}_{v\in c,n=1,\dots,N}$. By smearing all eigenvalues $\lambda_{n}(v)$, we have a set of intervals:
$I_{c}$ $=\bigcup_{v\in c,n=1,\dots,N}\Bigg{\{}\arg(\lambda_{n}(v)e^{i\delta\phi})\in[0,2\pi)\Bigg{|}$
$\hskip 30.0pt-\frac{\beta}{2N}<\delta\phi<\frac{\beta}{2N}\Bigg{\}}.$ (15)
Here, $10<\beta<$ 1 is a constant smearing parameter ensuring that adjacent eigenvalues fall within the same smeared interval. For example. we can set as $\beta=1/2$. We select a $\theta$ from the set $[0,2\pi)\backslash I_{c}$ to serve as $\theta_{c}$ for cube $c$. (Refer to Fig. 1 (c) for visualization.) With $\theta$-gaps for all cubes in lattice $L$ thus defined, the gap parameters $\theta_{p}^{+}$ and $\theta_{p}^{-}$ for each plaquette $p$ are specified. For each corner vertex $v_{0},v_{1},v_{2},v_{3}$ of plaquette $p$, we define $\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{j})$ as a $N\times N_{q}$ matrix:
$\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{a})=\left(u_{n_{1}}(v_{a}),\dots,u_{n_{N_{q}}}(v_{a})\right),$ (16)
comprising $N_{q}$ eigenvectors, with eigenvalue angles satisfying $\min(\theta_{p}^{+},\theta_{p}^{-})<\arg\lambda_{n_{1}}(v_{a})\leq\cdots\leq\arg\lambda_{n_{N_{q}}}(v_{a})<\max(\theta_{p}^{+},\theta_{p}^{-})$. The nonnegative integer $N_{q}$, indicating the count of eigenvalues between $e^{i\theta_{p}^{+}}$ and $e^{i\theta_{p}^{-}}$, should be in common for the four vertices $v_{0},v_{1},v_{2},v_{3}$ of plaquette $p$, assuming the lattice $L$ is fine enough. Then, the line integral $\oint_{\partial p}\alpha_{\theta_{p}^{-},\theta_{p}^{+}}$ can be approximated as
$\oint_{\partial p}\alpha_{\theta_{p}^{-},\theta_{p}^{+}}$ $\cong\mathrm{sgn}(\theta_{p}^{+}-\theta_{p}^{-})\times\mathrm{Arg}\,\det\left[\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{0})^{\dagger}\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{3})\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{3})^{\dagger}\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{2})\right.$
×$\hskip 50.0pt\left.\times\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{2})^{\dagger}\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{1})\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{1})^{\dagger}\gamma_{\{\theta_{p}^{+},\theta_{p}^{-}\}}(v_{0})\right]=:\Phi_{p},$ (17)
where $\mathrm{Arg}$ denotes the principal value ${-\pi<\mathrm{Arg}\ z<\pi}$. Consequently, we obtain a discrete formula for the winding number:
$W^{\rm dis}_{3}[g]=\frac{1}{2\pi}\sum_{p}\Phi_{p}\in\mathbb{Z},$ (18)
which relies solely on the diagonalization of matrix $g(v)$ at vertices $v\in L$ of the cubic lattice $L$ approximating the manifold $X$.
The discrete formula (18) is inherently quantized. To demonstrate this, we express $e^{2\pi iW^{\mathrm{dis}}_{3}[g]}$ as the product of edge contributions:
$e^{2\pi iW^{\mathrm{dis}}_{3}[g]}=\prod_{p}e^{i\Phi_{p}}$
$=\prod_{v_{a}v_{b}\in\{\mathrm{edges}\}}\exp\Bigg{[}i\arg\Big{\{}\det\left[\gamma_{\{\theta_{1},\theta_{0}\}}(v_{a})^{\dagger}\gamma_{\{\theta_{1},\theta_{0}\}}(v_{b})\right]^{\mathrm{sgn}(\theta_{1}-\theta_{0})}\det\left[\gamma_{\{\theta_{2},\theta_{1}\}}(v_{a})^{\dagger}\gamma_{\{\theta_{2},\theta_{1}\}}(v_{b})\right]^{\mathrm{sgn}(\theta_{2}-\theta_{1})}$
×$\quad\times\det\left[\gamma_{\{\theta_{3},\theta_{2}\}}(v_{a})^{\dagger}\gamma_{\{\theta_{3},\theta_{2}\}}(v_{b})\right]^{\mathrm{sgn}(\theta_{3}-\theta_{2})}\det\left[\gamma_{\{\theta_{0},\theta_{3}\}}(v_{a})^{\dagger}\gamma_{\{\theta_{0},\theta_{3}\}}(v_{b})\right]^{\mathrm{sgn}(\theta_{0}-\theta_{3})}\Big{\}}\Bigg{]}.$ (19)
Here, $v_{a}v_{b}$ denotes an individual edge, and $\theta_{0},\dots,\theta_{3}$ are the gap parameters for cubes adjacent to the edge $v_{a}v_{b}$, ordered counterclockwise from the vector $\overrightarrow{v_{b}v_{a}}$ pointing out of the page. See Fig. 2 (b). Regardless of the relative magnitudes of $\theta_{0},\theta_{1},\theta_{2}$, and $\theta_{3}$, each edge’s contribution in (19) cancels out exactly due to the property that for $\theta<\theta^{\prime}<\theta^{\prime\prime}$,
$\gamma_{\{\theta,\theta^{\prime\prime}\}}(v)=\left(\gamma_{\{\theta,\theta^{\prime}\}}(v_{a}),\gamma_{\{\theta^{\prime},\theta^{\prime\prime}\}}(v_{a})\right).$ (20)
Thus, $e^{2\pi iW^{\mathrm{dis}}_{3}[g]}=1$ is valid for any sufficiently fine discrete approximation of $X$.
Model calculation— Our formulation extends to computing the winding number for maps $g:X\to GL_{N}(\mathbb{C})$, where the target space consists of invertible matrices. Invertible matrices that cannot be diagonalized, known as ”exceptional points,” constitute a ring in the three-dimensional parameter space and are stable under minor perturbations. To circumvent these exceptional points, one can employ the singular value decomposition $g=U\Sigma V^{\dagger}$ to derive the unitary matrix $UV^{\dagger}$ at each vertex within the discretized parameter space. We verified our formula (18) with the model $g(k_{x},k_{y},k_{z})=t(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}+\sin k_{z}\sigma_{z})-i(m+\cos k_{x}+\cos k_{y}+\cos k_{z})\mathbf{1}_{2}$ on the three-torus $(k_{x},k_{y},k_{z})\in[-\pi,\pi]^{\times 3}$, employing a cubic lattice of $20\times 20\times 20$ mesh. Here, $\sigma_{\mu}\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$ denotes the Pauli matrices. We have checked that the winding number $W^{\rm dis}_{3}[g]$ equals $-2\mathrm{sgn}(t)$ for $|m|<1$, $\mathrm{sgn}(t)$ for $31<|m|<$ 3, and 00 for $|m|>3$, which is consistent with the direct calculation of the analytic form $W_{3}[g]$.
Summary— In this work, we presented a formulation for calculating the three-dimensional winding number $W_{3}[g]$ for smooth maps $g:X\to U(N)$, utilizing a discrete approximation of the manifold $X$. Our approach allows for the computation of $W_{3}[g]$ exclusively through the diagonalization of matrices $g(v)$ at a finite number of vertices, ensuring the result is explicitly quantized to integer values. Discrete formulations that explicitly quantify topological invariants are currently limited in scope. Examples such as instanton numbers represented by the second Chern numbers, higher-dimensional winding numbers, and degrees of maps to more general symmetric spaces have yet to be explored. We look forward to future studies shedding more light on these topics.
Acknowledgements.
We were supported by JST CREST Grant No. JPMJCR19T2, and JSPS KAKENHI Grant No. 22H05118 and 23H01097.