### v3

# A discrete formulation for three-dimensional winding number

Abstract
For a smooth map $g:X\to U(N)$, where $X$ is three-dimensional, oriented, and closed manifold, the winding number is defined as $W_{3}=\frac{1}{24\pi^{2}}\int_{X}\mathrm{Tr}\left[(g^{-1}dg)^{3}\right]$. We introduce a method to compute $W_{3}$ using a discrete approximation of $X$ that manifestly ensures the quantization of $W_{3}$.
preprint: YITP-24-14
Introduction— Consider a three-dimensional closed and oriented manifold $X$. For a smooth map $g:X\to U(N)$, with $U(N)$ representing the group of $N\times N$ unitary matrices, the winding number, an integer value, is defined by the following expression:
$W_{3}[g]$ $=\frac{1}{2\pi}\int_{X}H\in\mathbb{Z},$ (1)
$H$ $=\frac{1}{12\pi}\mathrm{Tr}\left[(g^{-1}\mathrm{d}g)^{3}\right].$ (2)
The winding number $W_{3}[g]$ is pivotal in various branches of physics, including topological band theory, where it acts as the topological invariant for three-dimensional superconductors with time-reversal symmetry [1, 2], and in non-Abelian (lattice) gauge theory, appears in instanton number calculations [3, 4]. Often in these applications, the function $g$ is defined only on a finite set of lattice points for numerical analysis. Therefore, an efficient numerical formulation with lattice approximation of manifolds is an important issue.
For the first Chern number $ch_{1}$ of a line bundle with connection, a sort of two-dimensional counterpart of the winding number, a well-established discrete formulation with evident quantization exists [5, 6]. Furthermore, discrete line bundles over finite simplicial complexes have been explored, especially concerning applications in computer graphics [7]. This paper develops a method for evaluating $W_{3}[g]$ via a discretized approximation of $X$, ensuring the result remains manifestly quantized, provided the approximation of $X$ is sufficiently refined relative to the spatial variation’s length scale.
Figure 1: (a) Illustration of a $\theta$-gap. (b) Projection onto the eigenspace between two $\theta$-gaps. (c) Smearing eigenvalues over a finite set of vertices. In each panel, the unit circle in the complex plane is depicted, with blue arcs representing the spectrum of the $U(N)$-valued matrix $g(x)$ within a local region. Red lines mark the locations of $\theta$-gaps.
Formulation— Given that $H$ is a closed three-form, it is locally exact, meaning that for a local patch, there exists a two-form $B$ such that $H=dB$. To construct $B$ explicitly, we introduce a gap condition for elements of $U(N)$. A matrix $g\in U(N)$ exhibits a $\theta$-gap if none of its eigenvalues are $e^{i\theta}$ for a given real number $\theta\in[0,2\pi)$ [8, 9], as illustrated in Fig. 1 (a). Furthermore, we define $\log_{\theta}z$ for nonzero complex number $z\in\mathbb{C}^{\times}$ as
$\log_{\theta}z=\log|z|+i\arg z,\quad\theta\leq\arg z<\theta+2\pi.$ (3)
For two distinct $\theta$-gaps, $\theta_{1}$ and $\theta_{2}$, the following relation holds:
$\log_{\theta_{1}}z-\log_{\theta_{2}}z$ =$=\begin{cases}2\pi i\times\mathrm{sgn}(\theta_{1}-\theta_{2})&(\min(\theta_{1},\theta_{2})<\arg z<\max(\theta_{1},\theta_{2})),\\ 0&(\mathrm{otherwise}).\end{cases}$ (4)
Consider $U\subset X$ as a three-dimensional subspace where $g(x)$ maintains a $\theta$-gap for $x\in U$. Let $\gamma(x)=(u_{1}(x),\dots,u_{N}(x))\in U(N)$ be a unitary matrix diagonalizing $g(x)$, i.e., $g(x)=\gamma(x)\Lambda(x)\gamma(x)^{\dagger}$ with $\Lambda(x)=\mathrm{diag}(\lambda_{1},\dots,\lambda_{N})$, where $\lambda_{n}\in U(1)$ for $n=1,\dots,N$. The exact form $B_{\theta}$ is given by [10]
$B_{\theta}$ $=Q+R_{\theta},$ (5)
$Q$ $=\frac{1}{4\pi}\mathrm{Tr}[\gamma^{-1}d\gamma\Lambda\gamma^{-1}d\gamma\Lambda^{-1}],$ (6)
$R_{\theta}$ $=\frac{1}{2\pi}\mathrm{Tr}[\log_{\theta}\Lambda(\gamma^{-1}d\gamma)^{2}].$ (7)
Note that while $Q$ is independent of $\theta$, but $R_{\theta}$ does. It is evident that $X$ can be covered by $\{U_{i}\}_{i}$ such that in each patch $U_{i}$, $g(x)$ exhibits a specific $\theta$-gap $\theta_{i}$ for $x\in U_{i}$.
The unitary matrix $\gamma$ is not unique because it can be transformed as $\gamma\mapsto\gamma W$, where $W\in U(N)$ commutes with $\Lambda$, i.e., $W\Lambda W^{-1}=\Lambda$. However, this ambiguity does not affect $Q$ and $R_{\theta}$: To clarify, consider the $N$ eigenvalues of $\Lambda$ grouped into blocks of $|I|$ degenerate eigenvalues, each associated with a common eigenvalue $\lambda_{I}$, so that $\Lambda$ can be expressed as $\Lambda=\bigoplus_{I}\lambda_{I}\mathbf{1}_{|I|}$. Introduce a block matrix notation $A_{IJ}=(u_{i}^{\dagger}du_{j})_{i\in I,j\in J}$. The transformation matrix $W$ is expressed as $W=\bigoplus_{I}W_{I}$ as well, with $W_{I}\in U(|I|)$, modifying $A_{IJ}$ to $A_{IJ}=W_{I}^{\dagger}A_{IJ}W_{J}+\delta_{IJ}W_{I}^{-1}dW_{I}$. Consequently, $Q$ and $R_{\theta}$ can be represented as:
$Q$ $=\frac{1}{4\pi}\sum_{I,J;I\neq J}\mathrm{Tr}_{I}[A_{IJ}A_{JI}]\lambda_{J}\lambda_{I}^{-1},$ (8)
$R_{\theta}$ $=\frac{1}{2\pi}\sum_{I,J;I\neq J}\mathrm{Tr}_{I}[A_{IJ}A_{JI}]\log_{\theta}\lambda_{I},$ (9)
where $\mathrm{Tr}_{I}$ denotes the trace over indices $i\in I$. In the summation $\sum_{I,J}$, terms with $I=J$ can be excluded due to $\mathrm{Tr}_{I}[(A_{II})^{2}]=0$. This demonstrates the invariance of $Q$ and $R_{\theta}$ under the transformation $\gamma\mapsto\gamma W$.
Another noteworthy aspect is that the difference in $B_{\theta}$ between two $\theta$-gaps is a total derivative. For $0\leq\theta_{1},\theta_{2}<2\pi$, and using (4), it follows that:
$B_{\theta_{1}}-B_{\theta_{2}}=d\alpha_{\theta_{1},\theta_{2}},$ (10)
where
$\alpha_{\theta_{1},\theta_{2}}$ $=-i\ \mathrm{sgn}(\theta_{1}-\theta_{2})\mathrm{Tr}[P_{\theta_{1},\theta_{2}}\gamma^{-1}d\gamma]$
$=-i\ \mathrm{sgn}(\theta_{1}-\theta_{2})$
×$\quad\times\sum_{n;\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})}u_{n}^{\dagger}du_{n},$ (11)
with
$P_{\theta_{1},\theta_{2}}=\sum_{n;\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})}u_{n}u_{n}^{\dagger},$ (12)
the orthogonal projection onto the eigenspace for eigenvalues that fulfill $\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})$. In cases where $\theta_{1}=\theta_{2}$ or no eigenvectors meet the condition $\min(\theta_{1},\theta_{2})<\arg\lambda_{n}<\max(\theta_{1},\theta_{2})$, $\alpha_{\theta_{1},\theta_{2}}$ is simply null.
Now, we express the winding number $W_{3}[g]$ as a sum of line integrals, utilizing a cubic lattice approximation $L$ of the manifold $X$. (Any simplicial decomposition is equally valid.) For each cube $c$ of the lattice $L$, we select $\theta_{c}\in[0,2\pi)$ such that $g_{x}$ for $x\in c$ exhibits a $\theta$-gap of $\theta_{c}$. Thus, $W_{3}[g]$ can be reformulated as a sum of integrals over all plaquettes:
[…]
$=\frac{1}{2\pi}\sum_{c}\int_{\partial c}B_{\theta_{c}}$
$=\frac{1}{2\pi}\sum_{p}\int_{p}(B_{\theta_{p}^{-}}-B_{\theta_{p}^{+}}).$ (13)
Here, $\sum_{p}$ runs over all plaquettes $p$ in the lattice $L$, with each $p$ being oriented. The gap parameters $\theta_{p}^{+}$ and $\theta_{p}^{-}$ correspond to the cubes adjacent to plaquette $p$, in directions parallel and antiparallel to $p$’s normal vector, respectively, as depicted in Fig. 2 (a). This expression further simplifies to a sum of line integrals:
[…]
$=\frac{1}{2\pi}\sum_{p}\oint_{\partial p}\alpha_{\theta_{p}^{-},\theta_{p}^{+}}.$ (14)
[…]
Figure 2: (a) A plaquette $p$ within the cubic lattice, showing $\theta_{p}^{+}$ and $\theta_{p}^{-}$ as the $\theta$-gaps of cubes adjacent to $p$, aligned parallel and anti-parallel to $p$’s normal vector, respectively. The vertices $v_{0},v_{1},v_{2},$ and $v_{3}$ are sequentially labeled around the perimeter of plaquette $p$. (b) An edge $v_{a}v_{b}$ of the cubic lattice, illustrating the $\theta$-gaps of cubes adjacent to the edge $v_{a}v_{b}$.
We claim that the winding number, as expressed in (14), can be approximately calculated using only the diagonalizing matrices $\gamma(v)$ at the vertices $v$ of the lattice $L$, while ensuring that $W_{3}[g]$ remains quantized. Diagonalizing $g(v)$ at vertices $v\in L$ yields the set of pairs of eigenvectors and eigenvalues $\{u_{n}(v),\lambda_{n}(v)\}_{n=1,\dots,N}$ for each vertex $v$. The eigenvalues $\lambda_{n}(v)$ are labeled such that their angles are in ascending order, satisfying
$0\leq\arg(\lambda_{1}(v))\leq\cdots\leq\arg(\lambda_{N}(v))<2\pi.$ (15)
This labeling prevents the sign factor that arises from reordering the eigenvectors over vertices when the Berry connection is evaluated as the inner product of eigenvectors among two adjacent vertices. Note also that eigenvalues $\lambda$ near 00 may reorder significantly under minor perturbations, yet this does not contribute to the discrete formula below. We introduce a notation for the set of eigenstates between two angles: For a given pair of gap parameters $0\leq\theta_{1},\theta_{2}<2\pi$, we define
$\gamma_{\theta_{1},\theta_{2}}(v):=\left(u_{n_{1}}(v),u_{n_{1}+1}(v),\dots,u_{n_{2}}(v)\right),$ (16)
where the eigenstates satisfy $\min(\theta_{1},\theta_{2})<\arg\left(\lambda_{n_{1}}(v)\right)\leq\cdots\leq\arg\left(\lambda_{n_{2}}(v)\right)<\max(\theta_{1},\theta_{2})$.
The gap parameter $\theta_{c}$ for each cube $c$ is numerically determined as follows: From the eight corner vertices of cube $c$, denoted as $v\in c$, we consider the set of $8N$ eigenvalues $\{\lambda_{n}(v)\}_{v\in c,n=1,\dots,N}$. By smearing all eigenvalues $\lambda_{n}(v)$, we have a set of intervals:
$I_{c}$ $=\bigcup_{v\in c,n=1,\dots,N}\Bigg{\{}\arg(\lambda_{n}(v)e^{i\delta\phi})\in[0,2\pi)\Bigg{|}$
$\hskip 30.0pt-\frac{\beta}{2N}<\delta\phi<\frac{\beta}{2N}\Bigg{\}}.$ (17)
Here, $10<\beta<$ 1 is a constant smearing parameter ensuring that near eigenvalues fall within the same smeared interval. For example. we can set as $\beta=1/2$. We select a $\theta$ from the complement $[0,2\pi)\backslash I_{c}$ to serve as $\theta_{c}$ for cube $c$. (See Fig. 1 (c) for visualization.) With $\theta$-gaps for all cubes in lattice $L$ thus defined, the gap parameters $\theta_{p}^{+}$ and $\theta_{p}^{-}$ for each plaquette $p$ are specified.
For the corner vertices $v_{0},v_{1},v_{2},v_{3}$ of a plaquette $p$, we define the eigenframes as
$\gamma_{\theta_{p}^{+},\theta_{p}^{-}}(v_{a})=\left(u_{n_{1}}(v_{a}),\dots,u_{n_{2}}(v_{a})\right)$ (18)
for $a=0,1,2,3$, comprising the $n_{2}-n_{1}$ eigenvectors with eigenvalues between $\theta_{p}^{+}$ and $\theta_{p}^{-}$. The nonnegative integer $n_{2}-n_{1}$, representing the count of eigenvalues between $\theta_{p}^{+}$ and $\theta_{p}^{-}$, should be consistent across the four vertices $v_{0},v_{1},v_{2},v_{3}$ of plaquette $p$, provided that the lattice $L$ is sufficiently fine. The line integral $\oint_{\partial p}\alpha_{\theta_{p}^{-},\theta_{p}^{+}}$ can then be approximated by the discrete Berry phase as
$\oint_{\partial p}\alpha_{\theta_{p}^{-},\theta_{p}^{+}}$ $\cong\mathrm{Arg}\Big{[}e^{iA_{v_{3}v_{0}\to p}}e^{iA_{v_{2}v_{3}\to p}}$
×$\times e^{iA_{v_{1}v_{2}\to p}}e^{iA_{v_{0}v_{1}\to p}}\Big{]}=:\Phi_{p},$ (19)
where
$e^{iA_{v_{a}v_{b}\to p}}=\det\left[\gamma_{\theta_{p}^{+},\theta_{p}^{-}}(v_{b})^{\dagger}\gamma_{\theta_{p}^{+},\theta_{p}^{-}}(v_{a})\right]^{\mathrm{sgn}(\theta_{p}^{+}-\theta_{p}^{-})}$ (20)
is the approximate edge $U(1)$ connection that contributes to the flux piercing the plaquette $p$. Here, $\mathrm{Arg}$ denotes the principal value, with $-\pi<\mathrm{Arg}\,z<\pi$.
Although the sum of fluxes $\frac{1}{2\pi}\sum_{p}\Phi_{p}$ provides a good approximation of the three-dimensional winding number $W_{3}[g]$, it is not generally quantized. To ensure that the sum is an integer, we express the sum of fluxes $\Phi_{p}$ modulo $2\pi$ in terms of edge contributions as follows:
$\sum_{p}\Phi_{p}$ $\equiv\sum_{v_{a}v_{b}\in\{\mathrm{edges}\}}\mathrm{Arg}\Bigg{[}e^{iA_{v_{a}v_{b}\to p_{01}}}e^{iA_{v_{a}v_{b}\to p_{12}}}e^{iA_{v_{a}v_{b}\to p_{23}}}e^{iA_{v_{a}v_{b}\to p_{30}}}\Bigg{]}$ (21)
$\equiv\sum_{v_{a}v_{b}\in\{\mathrm{edges}\}}\mathrm{Arg}\Bigg{[}\det\left[\gamma_{\theta_{0},\theta_{1}}(v_{b})^{\dagger}\gamma_{\theta_{0},\theta_{1}}(v_{a})\right]^{\mathrm{sgn}(\theta_{0}-\theta_{1})}\det\left[\gamma_{\theta_{1},\theta_{2}}(v_{b})^{\dagger}\gamma_{\theta_{1},\theta_{2}}(v_{a})\right]^{\mathrm{sgn}(\theta_{1}-\theta_{2})}$
×$\hskip 60.0pt\times\det\left[\gamma_{\theta_{2},\theta_{3}}(v_{b})^{\dagger}\gamma_{\theta_{2},\theta_{3}}(v_{a})\right]^{\mathrm{sgn}(\theta_{2}-\theta_{3})}\det\left[\gamma_{\theta_{3},\theta_{0}}(v_{b})^{\dagger}\gamma_{\theta_{3},\theta_{0}}(v_{a})\right]^{\mathrm{sgn}(\theta_{3}-\theta_{0})}\Bigg{]}\quad\bmod 2\pi.$ (22)
Here, $v_{a}v_{b}$ denotes an individual edge, $p_{01},p_{12},p_{23},p_{30}$ are the plaquettes adjacent to the edge $v_{a}v_{b}$, and $\theta_{0},\theta_{1},\theta_{2},\theta_{3}$ are the gap parameters such that the plaquette $p_{j,j+1}$ is adjacent to the two cubes with gap parameters $\theta_{j}$ and $\theta_{j+1}$. See Fig. 2 (b). An edge contribution ${\rm Arg}\left[e^{iA_{v_{a}v_{b}\to p_{01}}}e^{iA_{v_{a}v_{b}\to p_{12}}}e^{iA_{v_{a}v_{b}\to p_{23}}}e^{iA_{v_{a}v_{b}\to p_{30}}}\right]$ may not be zero. This occurs when an eigenframe $\gamma_{p_{j,j+1}}(v_{a})$ for a plaquette $p_{j,j+1}$ includes two or more other eigenframes.
For instance, consider the situation where $\theta_{0}<\theta_{1}<\theta_{2}=\theta_{3}$ and both the number of eigenvalues between $\theta_{0}$ and $\theta_{1}$, and $\theta_{1}$ and $\theta_{2}$, are nonzero. Then, $\gamma_{\theta_{3},\theta_{0}}(v)=\left(\gamma_{\theta_{0},\theta_{1}}(v),\gamma_{\theta_{1},\theta_{2}}(v)\right)$ holds for $v=v_{a},v_{b}$, and we have
$e^{iA_{v_{a}v_{b}\to p_{01}}}e^{iA_{v_{a}v_{b}\to p_{12}}}e^{iA_{v_{a}v_{b}\to p_{23}}}e^{iA_{v_{a}v_{b}\to p_{30}}}=\frac{\det\left[\begin{array}[]{cc}\gamma_{\theta_{0},\theta_{1}}(v_{b})^{\dagger}\gamma_{\theta_{0},\theta_{1}}(v_{a})&\gamma_{\theta_{0},\theta_{1}}(v_{b})^{\dagger}\gamma_{\theta_{1},\theta_{2}}(v_{a})\\ \gamma_{\theta_{1},\theta_{2}}(v_{b})^{\dagger}\gamma_{\theta_{0},\theta_{1}}(v_{a})&\gamma_{\theta_{1},\theta_{2}}(v_{b})^{\dagger}\gamma_{\theta_{1},\theta_{2}}(v_{a})\\ \end{array}\right]}{\det\left[\gamma_{\theta_{0},\theta_{1}}(v_{b})^{\dagger}\gamma_{\theta_{0},\theta_{1}}(v_{a})\right]\det\left[\gamma_{\theta_{1},\theta_{2}}(v_{b})^{\dagger}\gamma_{\theta_{1},\theta_{2}}(v_{a})\right]}.$ (25)
This is not a real value when the off-diagonal elements $\gamma_{\theta_{0},\theta_{1}}(v_{b})^{\dagger}\gamma_{\theta_{1},\theta_{2}}(v_{a})$ and $\gamma_{\theta_{1},\theta_{2}}(v_{b})^{\dagger}\gamma_{\theta_{0},\theta_{1}}(v_{a})$ are nonzero.
To eliminate such off-diagonal elements, we proceed as follows. Let $\theta_{0}^{\uparrow}\leq\theta_{1}^{\uparrow}\leq\theta_{2}^{\uparrow}\leq\theta_{3}^{\uparrow}$ be the increasing rearrangement of four angles $\theta_{0},\theta_{1},\theta_{2},\theta_{3}$, and define
$e^{iA^{\uparrow}_{v_{a}v_{b}}(j,j+1)}=\det\left[\gamma_{\theta_{j}^{\uparrow},\theta_{j+1}^{\uparrow}}(v_{b})^{\dagger}\gamma_{\theta_{j}^{\uparrow},\theta_{j+1}^{\uparrow}}(v_{a})\right],\quad j=0,1,2.$ (26)
Then, redefine the $U(1)$ connection $A_{v_{a}v_{b}\to p_{k,k+1}}$ as the diagonal sum of the block $U(1)$ connections $A^{\uparrow}_{v_{a}v_{b}}(j,j+1)$s between $\theta_{k}$ and $\theta_{k+1}$ as in
$e^{i\tilde{A}_{v_{a}v_{b}\to p_{k,k+1}}}:=\left[\prod_{j\in\{0,1,2\}:\min(\theta_{k},\theta_{k+1})\leq\theta_{j}^{\uparrow}\leq\theta_{j+1}^{\uparrow}\leq\max(\theta_{k},\theta_{k+1})}e^{iA^{\uparrow}_{v_{a}v_{b}}(j,j+1)}\right]^{{\rm sgn\,}(\theta_{k}-\theta_{k+1})}$ (27)
for $k=0,1,2,3$. (Note that $k+1=0$ when $k+1=4$ is interpreted cyclically.) With this modified $U(1)$ connection $e^{i\tilde{A}_{v_{a}v_{b}\to p}}$ contributing to the plaquette $p$ from the edge $v_{a}v_{b}$, we define the modified flux over the plaquette $p$ as
$\tilde{\Phi}_{p}:=\mathrm{Arg}\Big{[}e^{i\tilde{A}_{v_{3}v_{0}\to p}}e^{i\tilde{A}_{v_{2}v_{3}\to p}}e^{i\tilde{A}_{v_{1}v_{2}\to p}}e^{i\tilde{A}_{v_{0}v_{1}\to p}}\Big{]}.$ (28)
This leads to the discretized formula
$W^{\rm dis}_{3}[g]=\frac{1}{2\pi}\sum_{p}\tilde{\Phi}_{p}.$ (29)
By construction, it is evident that $e^{2\pi iW^{\mathrm{dis}}_{3}[g]}=1$, which ensures that $W^{\rm dis}_{3}[g]$ is an integer.
Model calculation— Our formulation extends to computing the winding number for maps $g:X\to GL_{N}(\mathbb{C})$, where the target space consists of invertible matrices. Invertible matrices that cannot be diagonalized, known as "exceptional points," constitute a ring in the three-dimensional parameter space and are stable under minor perturbations. To circumvent these exceptional points, one can employ the singular value decomposition $g=U\Sigma V^{\dagger}$ to derive the unitary matrix $UV^{\dagger}$ at each vertex within the discretized parameter space. We verified our formula (29) with the model $g_{0}(k_{x},k_{y},k_{z})=t(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}+\sin k_{z}\sigma_{z})-i(m+\cos k_{x}+\cos k_{y}+\cos k_{z})\mathbf{1}_{2}$ on the three-torus $(k_{x},k_{y},k_{z})\in[-\pi,\pi]^{\times 3}$, employing a cubic lattice of $20\times 20\times 20$ mesh. Here, $\sigma_{\mu}\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$ denotes the Pauli matrices. We have checked that the winding number $W^{\rm dis}_{3}[g_{0}]$ equals $-2\mathrm{sgn}(t)$ for $|m|<1$, $\mathrm{sgn}(t)$ for $31<|m|<$ 3, and 00 for $|m|>3$, which is consistent with the direct calculation of the analytic form $W_{3}[g_{0}]$. We have also verified that in the model given by $g_{1}=(g_{0}\oplus e^{\frac{\pi i}{2}}g_{0}\oplus U)+V$, where $g_{0}$ is as defined above, $U$ is a $2\times 2$ constant random unitary matrix, and $V$ is a $6\times 6$ constant perturbation, the lattice formula using $\Phi_{p}$ is not quantized, whereas the formula using $\tilde{\Phi}_{p}$ is quantized. A sample Mathematica code is available as an ancillary file on arXiv.
Summary— In this work, we presented a formulation for calculating the three-dimensional winding number $W_{3}[g]$ for smooth maps $g:X\to U(N)$, utilizing a discrete approximation of the manifold $X$. Our approach allows for the computation of $W_{3}[g]$ exclusively through the diagonalization of matrices $g(v)$ at a finite number of vertices, ensuring the result is explicitly quantized to integer values. Discrete formulations that explicitly quantify topological invariants are currently limited in scope. Examples such as instanton numbers represented by the second Chern numbers, higher-dimensional winding numbers, and degrees of maps to more general symmetric spaces have yet to be explored. We look forward to future studies shedding more light on these topics.
Acknowledgements.
I thank Tsuneya Yoshida for pointing out the error in the v1 version regarding the quantization of the lattice formula for the three-dimensional winding number. This work was supported by JST CREST Grant No. JPMJCR19T2, and JSPS KAKENHI Grant Nos. 22H05118 and 23H01097.